hamidiso23:

1. Central Angle: ?? 2. Arc Length of One Slice: ?? 3. Square Inches of Two Slices: ??

hamidiso23:

1 attachment
hamidiso23:

round to nearest tenth

hamidiso23:

@snowflake0531

hamidiso23:

@tranquility

crispyrat:

let me try and help Ok if the angle for the circle is 360 and then it is split into 8 parts than that means that the central angle is 360/8 which is?

hamidiso23:

45

crispyrat:

Ok then the arc length is a fraction of the circle's total circumference so you would after you find the central angle find the circumference of the whole thing and then multiply by 45/360=1/8

crispyrat:

C = 2*pi*r<--circumfrence area by the way

hamidiso23:

yup

crispyrat:

then you find the area of the whole pizza and then multiply it by 45*2/360=90/360=1/4

crispyrat:

@AZ is this right?

snowflake0531:

Arc length formula \(2 \pi r(\frac{\theta}{360})\) \(2\pi(6)(\frac{45}{360})\)

snowflake0531:

c. \(\frac{1}{2}{\pi r^2}\) \(\frac{1}{2}\pi (6^2))\)

hamidiso23:

so whats the answer

hamidiso23:

1/4

snowflake0531:

@crispyrat wrote:
let me try and help Ok if the angle for the circle is 360 and then it is split into 8 parts than that means that the central angle is 360/8 which is?
@hamidiso23 wrote:
45

hamidiso23:

fr

hamidiso23:

45 is the answer

snowflake0531:

for the first one yes

hamidiso23:

how do we get #2

snowflake0531:

@snowflake0531 wrote:
Arc length formula \(2 \pi r(\frac{\theta}{360})\) \(2\pi(6)(\frac{45}{360})\)

hamidiso23:

yup

hamidiso23:

2 times pie?

snowflake0531:

yes

hamidiso23:

6.28318530718

hamidiso23:

right

hamidiso23:

then do we multiplay that by 6

snowflake0531:

... type it into your browser

1 attachment
hamidiso23:

4.71

hamidiso23:

is the answer for #2

snowflake0531:

'round to the nearest tenth'

hamidiso23:

4.7

snowflake0531:

yes

hamidiso23:

now for # 3

snowflake0531:

@snowflake0531 wrote:
c. \(\frac{1}{2}{\pi r^2}\) \(\frac{1}{2}\pi (6^2))\)

hamidiso23:

1/2 times pie?

snowflake0531:

JuSt PuT tHe WhOlE tHiNg InTo YoUr BrOwSeR iTs EaSiEr

hamidiso23:

56.5486677644

hamidiso23:

mb

hamidiso23:

56.5

hamidiso23:

or 57

snowflake0531:

@hamidiso23 wrote:
56.5
yes

hamidiso23:

or 57?

hamidiso23:

56.5?

snowflake0531:

@snowflake0531 wrote:
@hamidiso23 wrote:
56.5
yes

hamidiso23:

ok tnx

SadHoodie:

u got it?